# Cross-domain EEG-based Emotion Recognition with Contrastive Learning

### v1

@@ Abstract
Electroencephalogram (EEG)-based emotion recognition is vital for affective computing but faces challenges in feature utilization and cross-domain generalization. This work introduces EmotionCLIP, which reformulates recognition as an EEG-text matching task within the CLIP framework. A tailored backbone, SST-LegoViT, captures spatial, spectral, and temporal features using multi-scale convolution and Transformer modules. Experiments on SEED and SEED-IV datasets show superior cross-subject accuracies of 88.69% and 73.50%, and cross-time accuracies of 88.46% and 77.54%, outperforming existing models. Results demonstrate the effectiveness of multimodal contrastive learning for robust EEG emotion recognition.
@@ 1 Introduction
Emotion recognition plays a vital role in mental health monitoring, early diagnosis of brain disorders, and enhancing human-computer interaction [goleman2020emotional]. Among different modalities, electroencephalogram (EEG)-based methods have become mainstream due to their non-invasiveness, high temporal resolution, and resistance to deliberate manipulation. Compared with facial expressions, speech, or gestures, EEG signals directly reflect brain activity associated with emotions, enabling reliable applications in healthcare, intelligent driving, and interactive systems.
Early methods relied on handcrafted features with support vector machine(SVM) [wang2011eeg] or k-nearest neighbor classification(KNN) [bahari2013eeg], which required expert knowledge and were sensitive to noise. In contrast, deep learning enables end-to-end feature learning and has significantly advanced EEG-based emotion recognition. Convolutional neural networks (CNNs) [deng2021sfe] and recurrent neural network (RNNs) [zhang2018strnn] capture local dependencies, and graph convolutional networks (GCNNs) [song2018dgcnn] and spatial-temporal RNNs [zhang2018spatial] model EEG’s topology and dynamics. Recently, Transformer-based models [zhai2025one], such as compact convolutional Transformers [song2022eegconformer] and multidimensional attention models [xu2023amdet], further improved performance by capturing spatio-temporal-frequency features.
Since EEG signals exhibit large inter-subject variability and non-stationarity, EEG-based emotion recognition undergoes performance degradation across subjects and time [zheng2016personalizing]. Domain adaptation and generalization methods [siddhad2024enhancing][zhao2025federated] aim to align source and target distributions or extract invariant features, but their effectiveness is limited by substantial cross-subject differences. Recently, contrastive learning has shown promise by enhancing robustness through similarity learning. For example, SimCLR-based methods [chen2020simple] and the CLISA framework [shen2023contrastive] improved inter-subject alignment by enforcing consistent representations across subjects exposed to the same stimuli. However, it is challenging to construct effective positive/negative pairs in contrastive learning methods.
To overcome these limitations, we propose EmotionCLIP, which reformulates EEG-based emotion recognition as an EEG-text matching task. EmotionCLIP integrates a pre-trained CLIP text encoder with a new EEG backbone, SST-LegoViT, that sequentially captures spatial, spectral, and temporal features via multi-scale convolution, Legoformer modules, and Transformer encoders. By aligning EEG features with semantic text embeddings, the model improves cross-subject and cross-time generalization. Experiments on SEED [seed] and SEED-IV [seediv] datasets show that EmotionCLIP achieves accuracies of 88.69% and 73.50% in cross-subject recognition, and 88.46% and 77.54% in cross-time recognition, outperforming state-of-the-art baselines such as MSFR-GCN [msfrgcn], MADA [MADA], and CLISA [shen2023contrastive].
@@
  • We introduce the CLIP framework into EEG emotion recognition, enabling cross-domain generalization through EEG-text matching.
  • We design the SST-LegoViT, a lightweight backbone that effectively learns EEG’s spatio-temporal-frequency information.
  • Extensive experiments demonstrate that our method achieves superior generalization and accuracy on benchmark datasets, establishing a new state of the art for cross-subject and cross-time recognition.
@@ 2.1 EmotionCLIP
Traditional cross-domain methods, such as domain adaptation [siddhad2024enhancing], aim to extract shared and invariant features directly from multi-source EEG data. However, this is challenging due to significant inter-subject variability and the non-stationary nature of EEG signals. Inspired by the success of the CLIP in multi-modal learning, we propose EmotionCLIP to reformulate the cross-domain emotion recognition task from a classification problem to an EEG-Text matching problem.
@@ 2.3 EEG Encoder
First, the signal is filtered into multiple frequency bands ( $\delta\,[1{-}4\,\text{Hz}]$, $\theta\,[4{-}8\,\text{Hz}]$, $\alpha\,[8{-}14\,\text{Hz}]$, $\beta\,[14{-}31\,\text{Hz}]$, $\gamma_{1}\,[31{-}\\ 51\,\text{Hz}]$ and $\gamma_{2}\,[51{-}75\,\text{Hz}]$). For each band, we extract Differential Entropy (DE) [de] and Power Spectral Density (PSD) [psd] features.
@@ 2.3.2 SST-LegoViT Architecture
Frequency-band Encoder (Legoformer): Different frequency-domain features (DE [de] and PSD [psd]) contribute differently to emotion recognition. To effectively and adaptively fuse them, we introduce the Legoformer module. As shown in Figure 4, This module consists of separate, parallel Transformer encoders for each feature type (one for DE, one for PSD). The outputs are then fused and fed into a decoder module, which uses cross-attention with the primary feature’s (DE) output as the key and value. This allows the model to leverage auxiliary features (PSD) while retaining focus on the most informative primary features (DE).
@@ 3.1 Datasets and Experimental Setup
All experiments were conducted on two publicly available EEG emotion datasets, SEED [seed] and SEED-IV [seediv]. To evaluate the model’s cross-domain generalization ability, two rigorous evaluation strategies were adopted: cross-subject validation using the leave-one-subject-out (LOSO) protocol, and cross-time validation across different sessions of the same subject. The model was trained with the AdamW optimizer, a batch size of 64, an initial learning rate of 0.0001 with a cosine decay schedule, and a weight decay of 0.003. Cross-entropy loss was employed, and an early stopping strategy was applied to terminate training if the validation accuracy did not improve for 50 consecutive epochs.

### v2

@@ Abstract
Electroencephalogram (EEG)-based emotion recognition is vital for affective computing but faces challenges in feature utilization and cross-domain generalization. This work introduces EmotionCLIP, which reformulates recognition as an EEG-text matching task within the CLIP framework. A tailored backbone, SST-LegoViT, captures spatial, spectral, and temporal features using multi-scale convolution and Transformer modules. Experiments on SEED and SEED-IV datasets show superior cross-subject accuracies of 88.69% and 73.50%, and cross-time accuracies of 88.46% and 77.54%, outperforming existing models. Results demonstrate the effectiveness of multimodal contrastive learning for robust EEG emotion recognition. The code is available at https://github.com/Departure2021/EmotionCLIP.
@@ 1 Introduction
Emotion recognition is vital in mental health monitoring, early diagnosis of brain disorders, and enhancing human-computer interaction. Among different modalities, EEG-based methods have become mainstream due to their non-invasiveness, high temporal resolution, and resistance to deliberate manipulation [8]. Compared with facial expressions, speech, or gestures, EEG signals directly reflect brain activity associated with emotions, enabling reliable applications in healthcare, intelligent driving, and interactive systems.
Early methods relied on handcrafted features with support vector machine [15] or k-nearest neighbor classification [1], which required expert knowledge and were sensitive to noise. Later, deep learning enables end-to-end feature learning and has significantly advanced EEG-based emotion recognition. Convolutional neural networks [3] captures local dependencies, and graph convolutional neural networks [13] and spatial-temporal recurrent neural networks [18] model EEG’s topology and dynamics. Recently, Transformer-based models [17], such as compact convolutional Transformers [14] and multidimensional attention models [16], further improved performance by capturing spatio-temporal-frequency features.
Since EEG signals exhibit large inter-subject variability and non-stationarity, EEG-based emotion recognition undergoes performance degradation across subjects and time [22]. Domain adaptation and generalization methods [12][19] aim to align source and target distributions or extract invariant features, but their effectiveness is limited by substantial cross-subject differences. Recently, contrastive learning has shown promise by enhancing robustness through similarity learning. For example, SimCLR-based methods [2] and the CLISA framework [11] improved inter-subject alignment by enforcing consistent representations across subjects exposed to the same stimuli. However, it is challenging to construct effective positive/negative pairs in contrastive learning methods.
To overcome these limitations, we propose EmotionCLIP, which reformulates EEG-based emotion recognition as an EEG-text matching task. EmotionCLIP integrates a pre-trained CLIP text encoder [9] with a new EEG backbone, SST-LegoViT, that sequentially captures spatial, spectral, and temporal features via multi-scale convolution, Legoformer modules, and Transformer encoders. By aligning EEG features with semantic text embeddings, the model improves cross-subject and cross-time generalization. Experiments on SEED [21] and SEED-IV [20] datasets show that EmotionCLIP achieves accuracies of 88.69% and 73.50% in cross-subject recognition, and 88.46% and 77.54% in cross-time recognition, outperforming state-of-the-art baselines such as MSFR-GCN [7], MADA [10], and CLISA [11].
@@
$\bullet$ We introduce the CLIP framework into EEG emotion recognition, enabling cross-domain generalization through EEG-text matching.
$\bullet$ We design the SST-LegoViT, a lightweight backbone that learns EEG’s spatio-temporal-frequency features.
$\bullet$ Extensive experiments demonstrate that our method achieves superior performance on benchmark datasets, establishing a new state of the art for cross-domain recognition.
@@ 2.1 EmotionCLIP
Traditional cross-domain methods, such as domain adaptation [12], aim to extract shared and invariant features directly from multi-source EEG data. However, this is challenging due to significant inter-subject variability and the non-stationary nature of EEG signals. Inspired by the success of the CLIP in multi-modal learning, we propose EmotionCLIP to reformulate the cross-domain emotion recognition task from a classification problem to an EEG-Text matching problem.
@@ 2.3 EEG Encoder
First, the signal is filtered into multiple frequency bands ( $\delta\,[1{-}4\,\text{Hz}]$, $\theta\,[4{-}8\,\text{Hz}]$, $\alpha\,[8{-}14\,\text{Hz}]$, $\beta\,[14{-}31\,\text{Hz}]$, $\gamma_{1}\,[31{-}\\ 51\,\text{Hz}]$ and $\gamma_{2}\,[51{-}75\,\text{Hz}]$). For each band, we extract Differential Entropy (DE) [4] and Power Spectral Density (PSD) [5] features.
@@ 2.3.2 SST-LegoViT Architecture
Frequency-band Encoder (Legoformer): Different frequency-domain features (DE [4] and PSD [5]) contribute differently to emotion recognition. To effectively and adaptively fuse them, we introduce the Legoformer module. As shown in Figure 4, This module consists of separate, parallel Transformer encoders for each feature type (one for DE, one for PSD). The outputs are then fused and fed into a decoder module, which uses cross-attention with the primary feature’s (DE) output as the key and value. This allows the model to leverage auxiliary features (PSD) while retaining focus on the most informative primary features (DE).
@@ 3.1 Datasets and Experimental Setup
All experiments were conducted on two publicly available EEG emotion datasets, SEED [21] and SEED-IV [20]. To evaluate the model’s cross-domain generalization ability, two rigorous evaluation strategies were adopted: cross-subject validation using the leave-one-subject-out (LOSO) protocol, and cross-time validation across different sessions of the same subject. The model was trained with the AdamW optimizer, a batch size of 64, an initial learning rate of 0.0001 with a cosine decay schedule, and a weight decay of 0.003. Cross-entropy loss was employed, and an early stopping strategy was applied to terminate training if the validation accuracy did not improve for 50 consecutive epochs.